OpenStudy (anonymous):

Describe the sequence of transformations from the related common function f(x)=x3 to g g(x)=2(x-8)^3

OpenStudy (perl):

h(x) = f(x-3) , first g(x) = 2*h(x)

OpenStudy (anonymous):

It's asking how it shifts and how it stretches or shrinks @perl

OpenStudy (anonymous):

It's a multiple choice answer